# A Causally Informed Pretraining Approach for Multimodal Foundation Models: Applications in Remote Sensing

### v3

@@ Abstract.
Self-supervised learning has emerged as a powerful paradigm for pretraining foundation models using large-scale data. Existing pretraining approaches predominantly rely on masked reconstruction or next-token prediction strategies, demonstrating strong performance across various downstream tasks, including geoscience applications. However, these approaches do not fully capture the causal interplay between different geospatial and environmental variables. To address this limitation, we propose Causally Informed Variable-Step Forecasting (CI-VSF), a novel pretraining task that models forecasting as a conditional generation task, where driver variables (e.g., weather) inform the prediction of response variables (e.g., satellite imagery). We demonstrate that pretraining in such a fashion leads to enhanced performance when finetuned on both prediction (e.g., crop mapping, missing image prediction, soil moisture estimation) and forecasting (e.g., future image forecasting, soil moisture forecasting) downstream tasks when compared to other pretraining approaches. While we use remote sensing as our main application to demonstrate the efficacy of our proposed pretraining strategy over existing paradigms, it is applicable to any domain that involves known causal relationships amongst a set of variables.
Pretraining Tasks, Causality, Foundation Model, Remote Sensing
@@ 1. Introduction
(c) Variable Step Forecasting (VSF)
@@
Figure 3. Our proposed novel Causally Informed Variable Step Forecasting (CI-VSF) task. Our pretraining task estimates a satellite image in the future (orange) using satellite imagery and weather context (yellow) and weather data up to that future date (green).
In recent years, self-supervised learning methods have emerged to utilize large amounts of data for pretraining model weights using specialized pretraining tasks (Gui et al., 2024; Jing and Tian, 2020). These models, known as foundation models, have demonstrated strong performance across various downstream tasks after fine-tuning(Awais et al., 2025; Liang et al., 2024; Zhou et al., 2024; Li et al., 2024). The predominant pretraining strategies come from either masked reconstruction (He et al., 2022; Devlin, 2018) or next-token prediction(Achiam et al., 2023). Beyond these, other approaches such as contrastive learning(Radford et al., 2021; Jaiswal et al., 2020), student-teacher models(Guo et al., 2024) have also been explored for pretraining.
@@
Masked reconstruction (MR)-based pretraining involves self-supervised tasks where a portion of the input data is hidden (masked), and the model is trained to reconstruct the original data. This style of pretraining has shown great progress in the vision domain (where parts of the image are masked(He et al., 2022)) and in the language domain (where words from sentence are masked(Devlin, 2018)). In the geoscience domain, masked reconstruction has been used extensively in the building of remote sensing foundation models (Jakubik et al., 2023; Cha et al., 2023; Cong et al., 2022; Tseng et al., 2023; Lu et al., 2024). In such models, parts of the satellite image are masked, and the objective is to reconstruct the entire satellite image (denoted by Figure 2(a)). Utility of finetuning such models has been demonstrated for numerous prediction based downstream tasks where the aim is to estimate or infer current or past variables using available data. Examples of such tasks include flood inundation mapping, wildfire scar mapping, cloud removal, urban semantic segmentation mapping, scene classification etc. (Jakubik et al., 2023; Cong et al., 2022; Sun et al., 2022). Researchers improved this pretraining task by adding multiple modalities (MM) to the reconstruction objective (shown in Figure 2(b)), where instead one modality, multiple masked modalities would be used to create a common embedding, from which all modalities would be reconstructed. Finetuning models trained with such a pretraining task has been shown to lead to improvements in scene classification and crop field classification (Nedungadi et al., 2024; Tseng et al., 2023).
@@
Next-token prediction-based pretraining involves self-supervised tasks where, given a sequence of data, the model is trained to predict the next or subsequent data points. This style of pretraining has been used in language modeling (where given a sequence of words the model learns to predict the next word(Touvron et al., 2023)) most notably GPT(Achiam et al., 2023). In the geoscience domain, next token prediction style of pretraining has been used in the climate/weather foundation models (Nguyen et al., 2023; Schmude et al., 2024; Pathak et al., 2022; Gao et al., 2022). Here, given a series of masked or unmasked climate variables, the objective is to estimate these climate variables at a random future timestamp (see Figure 2(c)). Models pretrained in this variable step forecasting (VSF) fashion have shown great performance in forecasting based downstream tasks, where the aim is to predict future variables based on past observations. Examples of such tasks include precipitation forecasting, hurricane trajectory tracking, sea surface temperature prediction.
In geosciences, there is often a complex interplay between different variables i.e., one set of variables (drivers) influence the other set of variables (response). For example, weather (driver) influences the vegetation growth (response) which can be observed by a satellite. In such scenarios none of the two pre-training paradigms (MR and VSF) fully utilize the relationship between these sets of variables for learning robust representations. For example, consider foundation models for remote sensing applications. Here, VSF based pre-training solely using satellite data is not adequate because change in satellite imagery in a region strongly depends on how weather interacts with that region. In other words, satellite imagery alone does not have enough information to estimate future states without weather. Furthermore, just using MR pretraining does not allow for capturing of temporal patterns in land cover change to be effectively included during pre-training.
In this paper, we propose a novel pre-training task for building foundation models that leverages the dependence (causality) between variables. This pre-training task, which we call Causally Informed Variable Step Forecasting (CI-VSF), models forecasting as a conditional generation task where conditioning is governed by the driver variables (e.g. weather) and the task is to predict the response variables (e.g. satellite images). While we use remote sensing as our main application to demonstrate the efficacy of this pretraining strategy over existing paradigms (MR and VSF), it is applicable to any domain that involves known causal relationships amongst a set of variables. Specifically, we propose forecasting of satellite imagery with weather context as a novel pre training task for building remote sensing foundation models (shown in Figure 3), i.e using a series of satellite imagery and weather data up till a future date, our task is to estimate the satellite image in the future. This method of pretraining captures the causal relationship between variables, and the resulting embeddings lead to better performance on both prediction and forecasting based downstream tasks compared to existing methods of pretraining.
To summarize, our key contributions are listed below:
$\bullet$ We propose Causally Informed Variable-Step Forecasting (CI-VSF), a pretraining objective that leverages causal-guided principles to train models while preserving the causal dependencies between modalities, unlike existing approaches such as standard multimodal masked reconstruction or variable-step forecasting.
$\bullet$ We show the efficacy of our pretraining task by training a foundation model in stages, beginning with multimodal masked reconstruction and ending with Causally Informed Variable-Step Forecasting.
$\bullet$ We demonstrate that embeddings when trained in our proposed stage-wise approach can outperform models trained with the baseline methods when fine-tuned for a range of spatiotemporal downstream applications such as soil moisture related tasks (prediction and forecasting), pixel wise crop mapping (prediction), and spectral imagery based tasks (prediction and forecasting)
$\bullet$ We release the code and model used to public (Link)
@@ 2. Related Work
Increased availability and ease of access to large scale satellite data has motivated the development of deep learning models that use this data to perform various geoscience tasks such as land cover mapping (Ghosh et al., 2021b; Kussul et al., 2017), wildfire mapping (Nayak et al., 2018; Seydi et al., 2022; Zhao et al., 2018), crop yield prediction (Kuwata et al., 2015; You et al., 2017), flood forecasting (Bentivoglio et al., 2022) etc. This has led to a large number of geoscience foundation models(Jakubik et al., 2023; Cha et al., 2023; Mall et al., 2023; Deng et al., 2024; Hong et al., 2024; Guo et al., 2024; Mai et al., 2022; Bastani et al., 2023; Lu et al., 2024; Schmude et al., 2024; Liu et al., 2024). Most of these foundation models can be placed into one of two groups based on the data they use : (1) weather-climate (Nguyen et al., 2023; Pathak et al., 2022) that are typically used for weather forecasting or climate modeling and (2) spectral data from remote sensing satellites (Cong et al., 2022; Jakubik et al., 2023; Tseng et al., 2023; Nedungadi et al., 2024; Liu et al., 2024) that are largely used for identifying land-use land-cover change dynamics. The most common pretraining task in geoscience has been reconstruction of spectral imagery. To enrich the embedding created, varying amounts of the input spectral image are masked, making the reconstruction task of the entire image harder leading to better embeddings (Cong et al., 2022). However, simple reconstruction embeddings capture just that particular image and might not be suited for downstream tasks that rely on multi temporal contexts such as crop mapping or land cover land use change. To solve this, previous works included multiple timestamps in their input, however some of these methods stacked these images together (Jakubik et al., 2023), thus removing the temporal aspect. However, some methods added a timestamp positional embedding so that the model has a sense of time (Khanna et al., 2023; Cong et al., 2022). This led to moderate success in handling downstream tasks that require multi temporal contexts. Another common pretraining task is forecasting of imagery. Typically, this pretraining task has been used in weather related foundation models and not in spectral imagery based foundation models. Foundation models created using masked forecasting has shown great success in weather related downstream tasks (Schmude et al., 2024; Pathak et al., 2022). To enrich the embeddings created, these works add variable future time forecasting, i.e vary the amount of time into the future the model needs to forecast, which was achieved by including an embedding for delta time (Nguyen et al., 2023). Other variants of foundation models include diffusion models that incorporate more information such as geographic location, time of year, country etc (Khanna et al., 2023). Other class of geoscience foundation models include contrastive learning based foundation models, where multiple modalities are aligned to form similar embeddings, which are then used for downstream tasks(Guo et al., 2024; Wang et al., 2024).
@@ 3. Architecture
Our architecture follows a heavy encoder and lightweight decoder format. Keeping our decoder lightweight forces richer embeddings from encoder, suitable for downstream tasks. As mentioned before, we incorporate multiple modalities (spectral imagery and weather) in pretraining our architecture. Figure 4 shows CI-VSF’s architecture. For the forecasting-based pre-training task, it is essential for architecture to capture spatiotemporal nature of satellite data, temporal nature of weather data, and most importantly, the causality.
Satellite image Encoder/Decoder: We use a shared Vision Transformer (ViT) to extract spatial features from spectral imagery across timestamps. ViT have been shown to be effective in the presence of high masking (He et al., 2022), even in geoscience contexts (Jakubik et al., 2023). The ViT converts each image into a patch grid of embeddings, incorporating patch positional information. This results in a series of spectral image embeddings on unmasked patches for each timestamp. Since our input is an image series, we propose using a shared ViT across timestamps, leading to a robust encoder, that is capable of embedding images from all timestamps.
Figure 4. Causally Informed Variable step Forecasting (CI-VSF) Architecture diagram
Weather Encoder Due to the coarse spatial resolution of weather data, typically for each image location, we have one weather data point value per timestamp. Thus we use a sequence-to-sequence Unidirectional LSTM to encode the weather data. The LSTM based approach showed higher accuracy over the transformer based approaches when trying to solely reconstruct masked weather. The LSTM generates weather embeddings for each timestamp, which are then subsampled to match input image timestamps, similar to WSTATT (Ravirathinam et al., 2024), called temporal embedding matching.
Timestamp/Delta Encoder For temporal information we incorporate day of year (DOY) using a shared linear layer with tanh activation, creating DOY embeddings for each timestamp. Additionally, we create another series that corresponds the number of days in between the images, i.e the delta in timestamps. We generate embeddings for these time delta between images using a separate linear layer. The DOY embeddings provide temporal context, while delta embeddings, shown useful in forecasting tasks (Nguyen et al., 2023), informs the model about forecast distance.
@@
For all but last embeddings in the series $Emb_{STW}$, we forecast the next time-step, i.e t1 of the input series would be used to forecast the image at t2, t2 would be used to forecast t3 and so on (In a similar fashion to the Causal Language Modeling pretraining task used in Large Language Models (Achiam et al., 2023)). For the last embedding $Emb^{t_{C}}_{:t_{C}}$ we forecast the K’th time-step, which is K-C days into the future. This K is a variable and is sampled for each training instance in a batch, thus the name variable-step forecasting (VSF). This method enables the model to predict future images based on current image embeddings and weather information up to the forecast date, acknowledging weather’s impact on land cover.
Decoder Following this stage, we repopulate these embeddings in their respective unmasked positions in the timestamps and zero out the masked patches to pass to the decoder, as done by most methods in transformer based autoencoder methods(Jakubik et al., 2023; He et al., 2022). We use a light-weight MLP decoder (similar to the one used in ViT) to ensure that the main focus of the model is to create strong encoder to capture the best information. Similar, to the encoder, we have a shared decoder that performs the operations on each timestamp using the same weights. The decoder maps the embeddings to the spectral image space and reshapes the output to match the required size of the spatiotemporal stack.

### v4

# Towards Knowledge Guided Pretraining Approaches for Multimodal Foundation Models: Applications in Remote Sensing

@@ Abstract.
Self-supervised learning has emerged as a powerful paradigm for pretraining foundation models using large-scale data. Existing pretraining approaches predominantly rely on masked reconstruction or next-token prediction strategies, demonstrating strong performance across various downstream tasks, including geoscience applications. However, these approaches do not fully capture the knowledge of causal interplay between different geospatial and environmental variables. To address this limitation, we propose Knowledge Guided Variable-Step Forecasting (KG-VSF), a novel pretraining task that models forecasting as a conditional generation task, where driver variables (e.g., weather) inform the prediction of response variables (e.g., satellite imagery). We demonstrate that pretraining in such a fashion leads to strong embeddings which give enhanced performance when finetuned on downstream tasks where capturing this causality matters such as pixel wise crop type mapping, soil moisture estimation and forecasting, missing image prediction, and future image forecasting when compared to finetuning embeddings from other standard pretraining approaches.
Pretraining Tasks, Knowledge Guided, Foundation Model, Remote Sensing
@@ 1. Introduction
(c) Single Modal Variable Step Forecasting (SM-VSF)
(d) Multi Modal Variable Step Forecasting (MM-VSF)
@@
Figure 3. Our proposed novel Knowledge Guided Variable Step Forecasting (KG-VSF) task. Our pretraining task estimates a satellite image in the future (orange) using satellite imagery and weather context (yellow) and weather data up to that future date (green).
Increased availability and ease of access to large scale satellite data has enabled the development of deep learning models for diverse geoscience applications, including land cover mapping (statt; kussul2017deep), wildfire mapping (nayak2018classifying; seydi2022burnt; zhao2018saliency), crop yield prediction (kuwata2015estimating; you2017deep), flood forecasting (bentivoglio2022deep) etc. This has also led to the emergence of a large number of geoscience foundation models(jakubik2023foundationmodelsgeneralistgeospatial; cha2023billion; mall2023remote; deng2024k2; hong2024spectralgpt; guo2024skysense; mai2022towards; bastani2023satlaspretrain; lu2024ai; schmude2024prithviwxcfoundationmodel; liu2024remoteclip).
Foundation models are a new class of machine learning models that are pre-trained using self-supervised learning approaches on large-scale datasets to generate high-quality embeddings(gui2024survey; jing2020self). These embeddings, after fine-tuning, have demonstrated strong performance across various downstream tasks (awais2025foundation; liang2024foundation; zhou2024comprehensive; li2024foundation). There are two broad categories of pretraining tasks that are used in most of the existing foundation models. The first category includes variants of masked reconstruction  (he2022masked; devlin2018bert) and next-token prediction(achiam2023gpt) that focus on reconstruction or sequence forecasting. The second category includes contrastive learning and distillation (chen2020simpleframeworkcontrastivelearningsimclr; caron2021emergingpropertiesselfsupervisedvisiondino), which focus on aligning embeddings across modalities. Our work in this paper builds on the first category by introducing a new pretraining task for use geospatial foundation models that leverages causal relationships among modalities; i.e. some variables (drivers) directly influence other variables (responses). The distinction from contrastive/distillation based approaches lies in the type of relationships being modeled, which will be revisited later for context.
@@
In masked reconstruction (MR)-based pretraining, portions of the input data are hidden (masked), and the model is trained to encode these masked input data into an embedding that is used to reconstruct the original (unmasked) data. This style of pretraining has shown great progress in the vision domain (where parts of the image are masked(he2022masked)) and in the language domain (where words from a sentence are masked(devlin2018bert)). In the geoscience domain, masked reconstruction has been used extensively in the building of remote sensing foundation models (jakubik2023foundationmodelsgeneralistgeospatial; cha2023billion; cong2022satmae; tseng2023lightweight; lu2024ai). In such models, parts of the satellite image are masked, and the pretraining objective is to reconstruct the entire satellite image (denoted by Figure 2(a)). Utility of embeddings produced via a single-modality masked reconstruction (SM-MR) pre-training objective has been demonstrated for numerous tasks where the aim is to estimate or infer current or past variables using available data. Examples of such tasks include flood inundation mapping, wildfire scar mapping, cloud removal, urban semantic segmentation mapping, scene classification etc. (jakubik2023foundationmodelsgeneralistgeospatial; cong2022satmae; sun2022ringmo). Variations of this pretraining task have been developed, where the pre-training task involves masking multiple modalities (MM) and reconstructing them (see Figure 2(b)), where instead of one modality, multiple masked modalities would be used. Embeddings produced by such a pre-training task (MM-MR - see Figure 2(b)) have been shown to lead to improvements in scene classification and crop field classification (nedungadi2024mmearth; tseng2023lightweight).
@@
In next-token prediction-based pretraining, the model learns to predict the next or subsequent data points in a given sequence. This style of pretraining has been used in language modeling (where given a sequence of words the model learns to predict the next word(touvron2023llama)) most notably GPT(achiam2023gpt). In the geoscience domain, next token prediction style of pretraining has been used in the climate/weather foundation models (nguyen2023climax; schmude2024prithviwxcfoundationmodel; pathak2022fourcastnet; gao2022earthformer). Here, given a series of masked or unmasked climate variables, the objective is to estimate these climate variables at a random future timestamp (see Figure 2(c)). This pretraining task has been augmented to include multiple modalities where instead of trying to forecast one variable, future values of multiple variables are estimated from a common embedding (see Figure 2(d)) Models pretrained in this variable step forecasting (VSF) fashion have shown great performance in tasks where the aim is to predict future variables based on past observations. Examples of such tasks include precipitation forecasting, hurricane trajectory tracking, and sea surface temperature prediction.
In geoscience applications, there is often a complex interplay between different variables i.e., one set of variables (drivers) influence the other set of variables (response). For example, weather (driver) influences the vegetation growth (response) which can be observed by a satellite. Note that the response observed by satellite in response to weather drivers will differ depending on the state/nature of the location. For example, grasslands may green up after rain, evergreen forests and urban areas may not change much in response to the rain, and different types of crops may be impacted by weather differently. Neither MR nor VSF attempt to capture such causal relationships between these sets of variables for learning robust embeddings. MR and VSF based pretraining primarily capture structure in the data in space and time, respectively, and do not explicitly leverage knowledge of causal relationship across modalities.
To address this gap, we propose Knowledge Guided Variable-Step Forecasting (KG-VSF), a novel pretraining objective that explicitly leverages known causal dependencies between modalities. In this framework, the pre-training task models forecasting as a conditional generation task where conditioning is governed by the driver variables (e.g. weather) and the task is to predict the response variables (e.g. satellite images). Specifically, given a series of weather data and satellite imagery (at irregular intervals) in the past and weather data up till a future date, our pretraining task is to estimate the satellite images in the future (see Figure 3). The embeddings produced in this pre-training task implicitly learn about the nature of the location (e.g., grass land vs forests vs specific crops) based only upon past weather and (irregularly available) satellite imagery, as this information determines differential response of weather on the future satellite imagery at that location. Note that past weather and satellite imagery are freely available on a global scale from government organizations (e.g., NASA, ESA), making this pre-training task useful in many geoscience domains.
Here, knowledge-guided refers to using domain knowledge of causal dependencies to define the pretraining objective itself. KG-VSF incorporates the understanding that weather variables act as drivers that influence surface conditions observed by satellites, and this directional relationship is explicitly represented in the learning task. This contrasts with prior multimodal fusion approaches, such as Tseng et al. (2023)(tseng2023lightweight), which combine multiple modalities without encoding directional influence.
Recent multimodal foundation models such as RemoteCLIP(liu2024remoteclip), SimCLR(chen2020simpleframeworkcontrastivelearningsimclr), DINO(caron2021emergingpropertiesselfsupervisedvisiondino), and SkySense(guo2024skysense) have demonstrated impressive success through contrastive or distillation-based objectives that align co-occurring representations across modalities. We discuss these approaches in detail in Section 2, as they represent an important and complementary direction. However, their objectives primarily capture semantic correlations between modalities rather than causal dependencies. In contrast, KG-VSF aims to learn causal dynamics, i.e., how one modality (e.g., weather) drives another (e.g., land surface state).
We evaluate the effectiveness of KG-VSF in settings where the impact of weather on land cover and its properties is visible via satellite imagery. For this we created two data sets (one global and the other for US crop lands), each incorporating Sentinel2 spectral data and ERA5 weather data. These data sets are used to pre-train models using our proposed KG-VSF method (Figure 3) and the 4 existing pretraining methods (Figure 2). We demonstrate that the embeddings produced from our proposed pretraining approach outperform embeddings produced from existing pretraining methods for the following downstream applications: soil moisture related tasks (prediction and forecasting), pixel wise crop type mapping (prediction), and spectral imagery based tasks (prediction and forecasting). We release the code and model used to public(Link).
The paper is organized as follows. Section 2 reviews related work and highlights how our approach differs from existing research. Section 3 describes the proposed deep learning architecture, and Section 4 details the pretraining task. Section 5 presents the data sources, masking scheme, and sampling strategy for the two datasets. Section 6 discusses the comparative pretraining frameworks, models, and implementation details. Section 7 focuses on the downstream task of crop type mapping, while Section 8 covers the downstream tasks of soil moisture prediction and forecasting. Section 9 examines the downstream tasks of missing-image prediction and future-image forecasting. Section 10 analyzes the quality of embeddings from different pretraining methods. Section 11 explores experiments to verify that KG-VSF captures causal relationships rather than mere correlations. Section 12 evaluates the scalability of the KG-VSF framework. Finally, Section 13 provides discussion and concluding remarks.
@@ 2. Related Work
Existing geoscience foundation models can be broadly grouped into two categories based on the data used: (1) weather-climate (nguyen2023climax; pathak2022fourcastnet) that are typically used for weather forecasting or climate modeling and (2) spectral data from remote sensing satellites (cong2022satmae; jakubik2023foundationmodelsgeneralistgeospatial; tseng2023lightweight; nedungadi2024mmearth; liu2024remoteclip) that are largely used for identifying land-use land-cover change dynamics.
The most common pretraining objective in remote sensing models is masked reconstruction of spectral imagery. To improve embedding richness, varying portions of input imagery are masked to make reconstruction harder, leading to stronger representations (cong2022satmae). However, embeddings trained this way capture primarily spatial features of individual images and may underperform on tasks that require temporal context, such as crop monitoring or land-cover change detection. To solve this, previous works included multiple timestamps in their input, however some of these methods stacked these images together (jakubik2023foundationmodelsgeneralistgeospatial), thus removing the temporal aspect. However, some methods added a timestamp positional embedding so that the model has a sense of time (khanna2023diffusionsat; cong2022satmae). This led to moderate success in handling downstream tasks that require multi temporal contexts.
Another pretraining strategy, forecasting-based pretraining, has been widely used in weather foundation models but less explored for spectral imagery. Models pretrained for masked forecasting have shown success in weather prediction tasks (schmude2024prithviwxcfoundationmodel; pathak2022fourcastnet). To enhance temporal generalization, variable-future-time forecasting has been introduced, where the forecast horizon varies and is encoded via a delta-time embedding [36]. Other variants include diffusion-based models, which integrate auxiliary contextual information such as geographic location, time of year, or region (nguyen2023climax).
A distinct line of work involves contrastive learning and distillation-based approaches such as RemoteCLIP(liu2024remoteclip), SimCLR(chen2020simpleframeworkcontrastivelearningsimclr), DINO(caron2021emergingpropertiesselfsupervisedvisiondino), and SkySense(guo2024skysense). Contrastive methods learn by comparing pairs of modalities and maximise similarity between their embeddings and also maximise dissimilarity between the embeddings of unrelated pairs. Distillation-based methods (e.g., DINO) instead train a student network to create an embedding that matches a slowly updated embedding produced by the teacher network by using different (augmented) views of the same input. While these approaches have been highly successful for cross-modal retrieval and semantic understanding, their pretraining objective focuses on co-occurrence alignment rather than modeling causal dependencies between different modalities. In contrast, as noted in the introduction, KG-VSF attempts to explicitly capture the causal impact of one modality over the other (e.g., “weather influences the state of the land cover”. Thus, the objective of contrastive/distillation models are complementary but orthogonal to the objective of KG-VSF, which is designed to capture causal knowledge flow within environmental processes, not semantic similarity across heterogeneous data sources.
@@ 3. Architecture
Our architecture follows a heavy encoder and lightweight decoder format. Keeping our decoder lightweight forces richer embeddings from encoder, suitable for downstream tasks. As mentioned before, we incorporate multiple modalities (spectral imagery and weather) in pretraining our architecture. Figure 4 shows KG-VSF’s architecture. For the forecasting-based pre-training task, it is essential for architecture to capture spatiotemporal nature of satellite data, temporal nature of weather data, and most importantly, create strong embeddings that capture knowledge of causality.
Satellite image Encoder/Decoder: We use a shared Vision Transformer (ViT) to extract spatial features from spectral imagery across timestamps. ViT have been shown to be effective in the presence of high masking (he2022masked), even in geoscience contexts (jakubik2023foundationmodelsgeneralistgeospatial). The ViT converts each image into a patch grid of embeddings, incorporating patch positional information. This results in a series of spectral image embeddings on unmasked patches for each timestamp. Since our input is an image series, we propose using a shared ViT across timestamps, leading to a robust encoder, that is capable of embedding images from all timestamps.
Figure 4. Knowledge Guided Variable step Forecasting (KG-VSF) Architecture diagram
Weather Encoder Due to the coarse spatial resolution of weather data, typically for each image location, we have one weather data point value per timestamp. Thus we use a sequence-to-sequence Unidirectional LSTM to encode the weather data. The LSTM based approach showed higher accuracy over the transformer based approaches when trying to solely reconstruct masked weather. The LSTM generates weather embeddings for each timestamp, which are then subsampled to match input image timestamps, similar to WSTATT (wstatt), called temporal embedding matching.
Timestamp/Delta Encoder For temporal information we incorporate day of year (DOY) using a shared linear layer with tanh activation, creating DOY embeddings for each timestamp. Additionally, we create another series that corresponds the number of days in between the images, i.e the delta in timestamps. We generate embeddings for these time delta between images using a separate linear layer. The DOY embeddings provide temporal context, while delta embeddings, shown useful in forecasting tasks (nguyen2023climax), informs the model about forecast distance.
@@
For all but last embeddings in the series $Emb_{STW}$, we forecast the next time-step, i.e t1 of the input series would be used to forecast the image at t2, t2 would be used to forecast t3 and so on (In a similar fashion to the Causal Language Modeling pretraining task used in Large Language Models (achiam2023gpt)). For the last embedding $Emb^{t_{C}}_{:t_{C}}$ we forecast the K’th time-step, which is K-C days into the future. This K is a variable and is sampled for each training instance in a batch, thus the name variable-step forecasting (VSF). This method enables the model to predict future images based on current image embeddings and weather information up to the forecast date, acknowledging weather’s impact on land cover.
Decoder Following this stage, we repopulate these embeddings in their respective unmasked positions in the timestamps and zero out the masked patches to pass to the decoder, as done by most methods in transformer based autoencoder methods(jakubik2023foundationmodelsgeneralistgeospatial; he2022masked). We use a light-weight MLP decoder (similar to the one used in ViT) to ensure that the main focus of the model is to create strong encoder to capture the best information. Similar, to the encoder, we have a shared decoder that performs the operations on each timestamp using the same weights. The decoder maps the embeddings to the spectral image space and reshapes the output to match the required size of the spatiotemporal stack.